at 2:00 p.m. a runner heads north on a highway jogging at 10 mph. at 2:30 p.m. a driver heads north on the same highway to pick up the runner. if the car travels at 55 mph, how long will it take the driver to catch the runner?
Answer by Alan3354(69443)

at 2:00 p.m. a runner heads north on a highway jogging at 10mph. at 2:30 p.m. a driver heads north on the same highway to pick up the runner. if the car travels at 55 mph, how long will it take the driver to catch the runner?
------------------
In 30 min, the runner goes 5 miles.
The driver gains on the runner at 45 mph (55-10)
5/45 = 1/9 hour = 20/3 minutes
= 6 min 40 seconds